# Efficient Neural Common Neighbor for Temporal Graph Link Prediction

### v1

@@ Appendix C Proofs
(Ability of encoding $k$-hop event). Given a $k$-hop event $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$. If the node embedding of $u_{0}$ at time $t_{u_{0},u_{1}}$, can be formally derived by the encoding function $\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\})$, then the learning method is considered capable of encoding the $k$-hop event. The following outline the encoding capabilities of different learning schemes:
@@
We consider the encoding of an arbitrary monotonically increasing $k$-hop temporal event sequence within a memory-based approach. The induction principle is applied to demonstrate the capability of this approach. For the base case, $k=1$, the encoding has been shown to be feasible. Now, assume the proposition holds for a $k\textquoteright$ ’-hop event; that is, any $k\textquoteright$ ’-hop temporal sequence of monotonically increasing events can be encoded using a memory-based approach. This assumption implies that there exists an embedding function such that
[…]
@@
$=f_{\textit{emb}}^{-1}(\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{1,\ldots,k^{\prime}\},k^{\prime}\geq 1\})$
[…]
We consider a $k$-hop-subgraph-based approach for our analysis. It is evident that a $k$-hop subgraph encompasses any $k\textquoteright$ ’-hop events, where $k\textquoteright\leq k$. Furthermore, the aggregation methodology assimilates all nodes contained within the subgraph. Collectively, these observations substantiate the theorem in question.
We commence with a restatement of Theorem 5.5 for clarity:
Theorem C.2.
@@
In the context of $k$-hop-subgraph-based algorithms wherein $k=1$, an event $(u,v,t)$ is exploited once for every incident event within the neighborhood of vertices $u$ or $v$. Without loss of generality, we focus on all events within the $1$-hop-subgraph of vertex $u$. The aggregate count of events processed is given by $\sum_{i=1}^{d(u)}i=\Theta\left(d(u)^{2}\right)$, where $d(u)$ denotes the degree of vertex $u$. Consequently, the computational complexity is fundamentally proportional to $\sum_{u\in{\mathcal{N}}}d(u)^{2}$. Drawing on the results of de Caen [1998], the lower bound on the time complexity is established as $\Omega\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}\right)$, whereas the upper bound is determined as ${\mathcal{O}}\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|\right)$
In the context of $k$-hop-subgraph-based algorithms wherein $k=2$,, we adopt similar strategy where each event$(u,v,t)$ will only be utilized once another event within the subgraph of $u$ or $v$ is firstly considered. The total number of events can be formulated as $\sum_{u\in{\mathcal{N}}}d(u)\sum_{v\in{\mathcal{N}}_{u}}\sum_{w\in{\mathcal{N}}_{v}}d(w)$. Replacing $d(u)$ as $X_{i}$, $\sum_{v\in{\mathcal{N}}_{u}}\sum_{w\in{\mathcal{N}}_{v}}$ as $Y_{i}$, we reformulated is as $\sum_{i\in|{\mathcal{N}}|}X_{i}Y_{i}$, satisfying $\sum_{i\in|{\mathcal{N}}|}X_{i}^{2}=\sum_{u\in{\mathcal{N}}}d(u)^{2}$ and $\sum_{i\in|{\mathcal{N}}|}Y_{i}^{2}=\sum_{u\in{\mathcal{N}}}d(u)^{4}$. Following Cauchy inequality and conclusions of $\sum_{u\in{\mathcal{N}}}d(u)^{2}$, we got the the upper-bound time complexity is ${\mathcal{O}}\left((\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|)^{\frac{3}{2}}\right)$
We commence with a restatement of Theorem 5.6 for clarity:
Theorem C.3.
@@
We first give definitions of these structural features under temporal settings. Given two nodes $u$ and $v$, the structural features before time $t$ are defined as follows:
@@
Given a node $u$, the degree of node $u$ is the number of events $e$ with an endpoint at node $u$. Without loss of generality (WLOG), we consider node $u$ as the source node, and the events are $\{(u,v_{i},t_{i})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$. Each time a new event is given, the embedding of node $u$ is updated by
@@
Thus, the neural common neighbor $TNCN_{1}(u,v)=\mathop{\oplus}_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}emb_{w}^{t}$ can express Equation 9.

### v2

@@ Appendix I Proofs
(Theorem 4.4. Ability of encoding $k$-hop event). Given a $k$-hop event $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$. If the node embedding of $u_{0}$ at time $t_{u_{0},u_{1}}$, can be formally derived by the encoding function $\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\})$, then the learning method is considered capable of encoding the $k$-hop event. The following outline the encoding capabilities of different learning schemes:
@@
We consider the encoding of an arbitrary monotonically increasing $k$-hop temporal event sequence within a memory-based approach. The induction principle is applied to demonstrate the capability of this approach. For the base case, $k=1$, the encoding has been shown to be feasible. Now, assume the proposition holds for a $k$’-hop event; that is, any $k$’-hop temporal sequence of monotonically increasing events can be encoded using a memory-based approach. This assumption implies that there exists an embedding function such that
$Emb(u_{0},t_{u_{0},u_{1}})=\textit{Enc}({(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in{0,\ldots,k’-1}}),$ (7)
for all event sequences with $k^{\prime}$ hops, where $k’\geq 1$. Given an arbitrary k+1k’+1-hop event, which can be partitioned into an initial event $(u_{0},u_{1},t_{u_{0},u_{1}})$ and a subsequent $k$’-hop sequence. The existence of an encoding function for the $k$’-hop sequence assures that
@@
$=f_{\textit{emb}}^{-1}(\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})$
$\mid i\in\{1,\ldots,k^{\prime}\},k^{\prime}\geq 1\})$
Subsequently, it is demonstrated that $Emb(u_{0},t_{u_{0},u_{1}})$ provides an encoding for both the initial event and the $k$’-hop sequence, thereby affirming its efficacy in encoding the entire k+1k’+1-hop event. This concludes the inductive step and substantiates the inductive argument.
We consider a $k$-hop-subgraph-based approach for our analysis. It is evident that a $k$-hop subgraph encompasses any $k$’-hop events, where $k’\leq k$. Furthermore, the aggregation methodology assimilates all nodes contained within the subgraph. Collectively, these observations substantiate the theorem in question.
Theorem I.2.
(Theorem 4.7. Learning method time complexity). Denote the time complexity of a learning method as a function of the total number of events processed during training. For a given graph $\mathcal{G}$} with the number of nodes designated as $|{\mathcal{N}}|$ and the number of edges as $|{\mathcal{E}}|$, the following assertions hold:
@@
In the context of $k$-hop-subgraph-based algorithms wherein $k=1$, an event $(u,v,t)$ is exploited once for every incident event within the neighborhood of vertices $u$ or $v$. Without loss of generality, we focus on all events within the $1$-hop-subgraph of vertex $u$. The aggregate count of events processed is given by $\sum_{i=1}^{d(u)}i=\Theta\left(d(u)^{2}\right)$, where $d(u)$ denotes the degree of vertex $u$. Consequently, the computational complexity is fundamentally proportional to $\sum_{u\in{\mathcal{N}}}d(u)^{2}$. Drawing on the results of DECAEN1998245, the lower bound on the time complexity is established as $\Omega\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}\right)$, whereas the upper bound is determined as ${\mathcal{O}}\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|\right)$
In the context of $k$-hop-subgraph-based algorithms wherein $k=2$, we adopt similar strategy where each event(u,v,t)(u,v,t) will only be utilized once another event within the subgraph of $u$ or $v$ is firstly considered. The total number of events can be formulated as $\sum_{u\in{\mathcal{N}}}d(u)\sum_{v\in{\mathcal{N}}_{u}}\sum_{w\in{\mathcal{N}}_{v}}d(w)$. Replacing $d(u)$ as $X_{i}$, $\sum_{v\in{\mathcal{N}}_{u}}\sum_{w\in{\mathcal{N}}_{v}}$ as $Y_{i}$, we reformulated is as $\sum_{i\in|{\mathcal{N}}|}X_{i}Y_{i}$, satisfying $\sum_{i\in|{\mathcal{N}}|}X_{i}^{2}=\sum_{u\in{\mathcal{N}}}d(u)^{2}$ and $\sum_{i\in|{\mathcal{N}}|}Y_{i}^{2}=\sum_{u\in{\mathcal{N}}}d(u)^{4}$. Following Cauchy inequality and conclusions of $\sum_{u\in{\mathcal{N}}}d(u)^{2}$, we got the the upper-bound time complexity is ${\mathcal{O}}\left((\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|)^{\frac{3}{2}}\right)$
Theorem I.3.
(Theorem 4.6. Expressivity of TNCN)
  1. 1.
@@
  2. 2.
    TNCN is strictly more expressive than Jodie with the same dimension of time encoding, DyRep with the same aggregation function, TGAT with the same attention layers and neighbors, and TGN under identical condition for all module choices.
Proof.
(1) We first give definitions of these structural features under temporal settings. Given two nodes $u$ and $v$, the structural features before time $t$ are defined as follows:
@@
Given a node $u$, the degree of node $u$ is the number of events $e$ with an endpoint at node $u$. Without loss of generality (W.L.O.G.), we consider node $u$ as the source node, and the events are $\{(u,v_{i},t_{i})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$. Each time a new event is given, the embedding of node $u$ is updated by
@@
Thus, the neural common neighbor
$TNCN_{1}(u,v)=\mathop{\oplus}_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}emb_{w}^{t}$ can express Equation 9.
@@
(2) We first prove that TNCN is strictly more expressive than TGN. As TNCN possesses the same memory-based framework as TGN to derive the $emb_{u}^{t}$, TNCN at least has the same expressivity as TGN with identical module choices. Additionally, from part (1) of this theorem we know that TNCN can capture certain heuristics such as CN where TGN fails. Hence TNCN is strictly more expressive than TGN.
Since Jodie, DyRep and TGAT are specific instances of TGN with minor changes, we next shift our object to prove that TGN has no less expressivity than the three methods.
For TGAT, it retains only the Graph Attention mechanism for embedding functions, removing all memory and message modules present in TGN. Consequently, TGAT lacks the capacity to incorporate temporal memory updates and message passing, making it less expressive than TGN.
Jodie differs from TGN in the memory updater and embedding function. Specifically, it utilizes an RNN rather than a GRU, although GRU is a general variant of RNN. Additionally, Jodie employs the time embedding as $emb_{u}^{t}=(1+\Delta t\cdot w)\otimes mem_{u}^{t^{-}}$, where $w$ are learnable parameters. However, TGN’s Graph Attention uses not only the memory $mem_{u}^{t^{-}}$ but also the historical edge features $e_{u,v}^{t^{\prime}}$, and learns the coefficient via multi-head attention, thus achieving a higher expressivity.
DyRep also modifies the memory updater to an RNN. Meanwhile, it changes the embedding function to Identity and message function to Graph Attention. Although it just swaps these two modules, the expressivity is reduced in the temporal setting. When the model attempts to make its prediction, it first generates the embeddings with the embedding function. So naturally the identity function processes less information in the $emb_{u}^{t}$ than graph attention. After obtaining the prediction result, the new messages will be calculated by the message function and fed into memory updater, where the $emb_{u}^{t}$ can not get access. Hence TGN is more expressive than DyRep, especially when the recent information can make large benefits.
In conclusion, TNCN is strictly more expressive than Jodie, DyRep, TGAT and TGN under the same condition.